# Electron-phonon-dominated charge-density-wave fluctuations in TiSe2 accessed by ultrafast nonequilibrium dynamics

### v1

@@ Acknowledgements.
Acknowledgments We thank Claude Monney for useful discussions. We thank Nikita Fedorov, Romain Delos, Pierre Hericourt, Rodrigue Bouillaud, Laurent Merzeau, and Frank Blais for technical assistance and Aymen Mahmoudi, Joel Morf, and Frederic Chassot for their assistance during STM measurements. J.S., S.S. and J.M. would like to thank the QM4ST project with Reg. No. CZ.02.01.01/00/22_008/0004572, cofunded by the ERDF as part of the MŠMT. The research leading to these results has received funding from LASERLAB-EUROPE (grant agreement no. 871124, European Union’s Horizon 2020 research and innovation programme). We acknowledge the financial support of the IdEx University of Bordeaux / Grand Research Program "GPR LIGHT". We acknowledge support from ERC Starting Grant ERC-2022-STG No.101076639, Quantum Matter Bordeaux, AAP CNRS Tremplin and AAP SMR from Université de Bordeaux. SF acknowledges funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie 2024 Postdoctoral Fellowship No 101198277 (TopQMat). JS acknowledges funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101209345—ART.QM funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or European Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work was funded by the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat (EXC 2147, project-id 390858490) and by the DFG through SFB1170 “Tocotronics” and directly via RE 1469/13-2. N.G.E., and D.N. acknowledge financial support from the project "Podizanje znanstvene izvrsnosti Centra za napredne laserske tehnike (CALTboost)" financed by the European Union through the National Recovery and Resilience Plan 2021-2026 (NRPP).References
- Ishioka et al. (2010) J. Ishioka, Y. H. Liu, K. Shimatake, T. Kurosawa, K. Ichimura, Y. Toda, M. Oda, and S. Tanda, Phys. Rev. Lett. 105, 176401 (2010).
- Zhang et al. (2022) R. Zhang, W. Ruan, J. Yu, L. Gao, H. Berger, L. Forró, K. Watanabe, T. Taniguchi, A. Ranjbar, R. V. Belosludov, T. D. Kühne, M. S. Bahramy, and X. Xi, Phys. Rev. B 105, 085409 (2022).
- Xu et al. (2020) S.-Y. Xu, Q. Ma, Y. Gao, A. Kogar, A. Zong, A. M. Mier Valdivia, T. H. Dinh, S.-M. Huang, B. Singh, C.-H. Hsu, T.-R. Chang, J. P. C. Ruff, K. Watanabe, T. Taniguchi, H. Lin, G. Karapetrov, D. Xiao, P. Jarillo-Herrero, and N. Gedik, Nature 578, 545–549 (2020).
- Morosan et al. (2006) E. Morosan, H. W. Zandbergen, B. S. Dennis, J. W. G. Bos, Y. Onose, T. Klimczuk, A. P. Ramirez, N. P. Ong, and R. J. Cava, Nature Physics 2, 544 (2006).
- Holt et al. (2001) M. Holt, P. Zschack, H. Hong, M. Y. Chou, and T.-C. Chiang, Phys. Rev. Lett. 86, 3799 (2001).
- Di Salvo et al. (1976) F. J. Di Salvo, D. E. Moncton, and J. V. Waszczak, Phys. Rev. B 14, 4321 (1976).
- Woo et al. (1976a) K. C. Woo, F. C. Brown, W. L. McMillan, R. J. Miller, M. J. Schaffman, and M. P. Sears, Phys. Rev. B 14, 3242 (1976a).
- Rossnagel (2011) K. Rossnagel, Journal of Physics: Condensed Matter 23, 213001 (2011).
- Kidd et al. (2002) T. E. Kidd, T. Miller, M. Y. Chou, and T.-C. Chiang, Phys. Rev. Lett. 88, 226402 (2002).
- Cercellier et al. (2007) H. Cercellier, C. Monney, F. Clerc, C. Battaglia, L. Despont, M. G. Garnier, H. Beck, P. Aebi, L. Patthey, H. Berger, and L. Forró, Phys. Rev. Lett. 99, 146403 (2007).
- Monney et al. (2009) C. Monney, H. Cercellier, F. Clerc, C. Battaglia, E. F. Schwier, C. Didiot, M. G. Garnier, H. Beck, P. Aebi, H. Berger, L. Forró, and L. Patthey, Phys. Rev. B 79, 045116 (2009).
- Monney et al. (2010) C. Monney, E. F. Schwier, M. G. Garnier, N. Mariotti, C. Didiot, H. Beck, P. Aebi, H. Cercellier, J. Marcus, C. Battaglia, H. Berger, and A. N. Titov, Phys. Rev. B 81, 155104 (2010).
- Monney et al. (2012) C. Monney, G. Monney, P. Aebi, and H. Beck, Phys. Rev. B 85, 235150 (2012).
- Kogar et al. (2017) A. Kogar, M. S. Rak, S. Vig, A. A. Husain, F. Flicker, Y. I. Joe, L. Venema, G. J. MacDougall, T. C. Chiang, E. Fradkin, J. van Wezel, and P. Abbamonte, Science 358, 1314 (2017).
- Watson et al. (2019) M. D. Watson, O. J. Clark, F. Mazzola, I. Marković, V. Sunko, T. K. Kim, K. Rossnagel, and P. D. C. King, Phys. Rev. Lett. 122, 076404 (2019).
- Jérome et al. (1967) D. Jérome, T. M. Rice, and W. Kohn, Phys. Rev. 158, 462 (1967).
- Hughes (1977) H. P. Hughes, Journal of Physics C: Solid State Physics 10, L319 (1977).
- Yoshida and Motizuki (1980) Y. Yoshida and K. Motizuki, Journal of the Physical Society of Japan 49, 898 (1980).
- Whangbo and Canadell (1992) M. H. Whangbo and E. Canadell, Journal of the American Chemical Society 114, 9587 (1992).
- Rossnagel et al. (2002) K. Rossnagel, L. Kipp, and M. Skibowski, Phys. Rev. B 65, 235101 (2002).
- Weber et al. (2011) F. Weber, S. Rosenkranz, J.-P. Castellan, R. Osborn, G. Karapetrov, R. Hott, R. Heid, K.-P. Bohnen, and A. Alatas, Phys. Rev. Lett. 107, 266401 (2011).
- Porer et al. (2014) M. Porer, U. Leierseder, J.-M. Ménard, H. Dachraoui, L. Mouchliadis, I. E. Perakis, U. Heinzmann, J. Demsar, K. Rossnagel, and R. Huber, Nature Materials 13, 857 (2014).
- Monney et al. (2016) C. Monney, M. Puppin, C. W. Nicholson, M. Hoesch, R. T. Chapman, E. Springate, H. Berger, A. Magrez, C. Cacho, R. Ernstorfer, and M. Wolf, Phys. Rev. B 94, 165165 (2016).
- Karam et al. (2018) T. E. Karam, J. Hu, and G. A. Blake, ACS Photonics 5, 1228 (2018).
- Hedayat et al. (2019) H. Hedayat, C. J. Sayers, D. Bugini, C. Dallera, D. Wolverson, T. Batten, S. Karbassi, S. Friedemann, G. Cerullo, J. van Wezel, S. R. Clark, E. Carpene, and E. Da Como, Phys. Rev. Research 1, 023029 (2019).
- Lian et al. (2020) C. Lian, S. J. Zhang, S. Q. Hu, M. X. Guan, and S. Meng, Nature Communications 11, 43 (2020).
- Otto et al. (2021) M. R. Otto, J.-H. Pöhls, L. P. R. de Cotret, M. J. Stern, M. Sutton, and B. J. Siwick, Science Advances 7, eabf2810 (2021).
- Cheng et al. (2022) Y. Cheng, A. Zong, J. Li, W. Xia, S. Duan, W. Zhao, Y. Li, F. Qi, J. Wu, L. Zhao, et al., Nature communications 13, 963 (2022).
- Heinrich et al. (2023) T. Heinrich, H.-T. Chang, S. Zayko, K. Rossnagel, M. Sivis, and C. Ropers, Phys. Rev. X 13, 021033 (2023).
- van Wezel et al. (2010) J. van Wezel, P. Nahai-Williamson, and S. S. Saxena, Phys. Rev. B 81, 165109 (2010).
- Rohwer et al. (2011) T. Rohwer, S. Hellmann, M. Wiesenmayer, C. Sohrt, A. Stange, B. Slomski, A. Carr, Y. Liu, L. Miaja-Avila, M. Kallaene, S. Mathias, L. Kipp, K. Rossnagel, and M. Bauer, Nature 471, 490 (2011).
- Hellmann et al. (2012) S. Hellmann, T. Rohwer, M. Kalläne, K. Hanff, C. Sohrt, A. Stange, A. Carr, M. Murnane, H. Kapteyn, L. Kipp, et al., Nature communications 3, 1069 (2012).
- Duan et al. (2021) S. Duan, Y. Cheng, W. Xia, Y. Yang, C. Xu, F. Qi, C. Huang, T. Tang, Y. Guo, W. Luo, D. Qian, D. Xiang, J. Zhang, and W. Zhang, Nature 595, 239 (2021).
- Woo et al. (1976b) K. C. Woo, F. C. Brown, W. L. McMillan, R. J. Miller, M. J. Schaffman, and M. P. Sears, Phys. Rev. B 14, 3242 (1976b).
- Miyahara et al. (1995) Y. Miyahara, H. Bando, and H. Ozaki, Journal of Physics: Condensed Matter 7, 2553 (1995).
- Chen et al. (2016) P. Chen, Y.-H. Chan, X.-Y. Fang, S.-K. Mo, Z. Hussain, A.-V. Fedorov, M. Chou, and T.-C. Chiang, Scientific reports 6, 37910 (2016).
- Mizukoshi et al. (2023) Y. Mizukoshi, T. Fukuda, Y. Komori, R. Ishikawa, K. Ueno, and M. Hase, Applied Physics Letters 122, 243101 (2023).
- Guo et al. (2025) X. Guo, A. Kogar, J. Henke, F. Flicker, F. de Juan, S. X. L. Sun, I. Khayr, Y. Peng, S. Lee, M. J. Krogstad, S. Rosenkranz, R. Osborn, J. P. C. Ruff, D. B. Lioi, G. Karapetrov, D. J. Campbell, J. Paglione, J. van Wezel, T. C. Chiang, and P. Abbamonte, arXiv , 2501.09968 (2025).
- Jaouen et al. (2019) T. Jaouen, B. Hildebrand, M.-L. Mottas, M. Di Giovannantonio, P. Ruffieux, M. Rumo, C. W. Nicholson, E. Razzoli, C. Barreteau, A. Ubaldini, E. Giannini, F. Vanini, H. Beck, C. Monney, and P. Aebi, Phys. Rev. B 100, 075152 (2019).
- Lian et al. (2019) C. Lian, Z. A. Ali, and B. M. Wong, Phys. Rev. B 100, 205423 (2019).
- Joe et al. (2014) Y. I. Joe, X. M. Chen, P. Ghaemi, K. D. Finkelstein, G. A. de la Peña, Y. Gan, J. C. T. Lee, S. Yuan, J. Geck, G. J. MacDougall, T. C. Chiang, S. L. Cooper, E. Fradkin, and P. Abbamonte, Nature Physics 10, 421 (2014).
- Hinlopen et al. (2024) R. D. H. Hinlopen, O. N. Moulding, W. R. Broad, J. Buhot, F. Bangma, A. McCollam, J. Ayres, C. J. Sayers, E. D. Como, F. Flicker, J. van Wezel, and S. Friedemann, Science Advances 10, eadl3921 (2024).
- Borisenko et al. (2008) S. V. Borisenko, A. A. Kordyuk, A. N. Yaresko, V. B. Zabolotnyy, D. S. Inosov, R. Schuster, B. Büchner, R. Weber, R. Follath, L. Patthey, and H. Berger, Phys. Rev. Lett. 100, 196402 (2008).
- Arpaia et al. (2019) R. Arpaia, S. Caprara, R. Fumagalli, G. D. Vecchi, Y. Y. Peng, E. Andersson, D. Betto, G. M. D. Luca, N. B. Brookes, F. Lombardi, M. Salluzzo, L. Braicovich, C. D. Castro, M. Grilli, and G. Ghiringhelli, Science 365, 906 (2019).
- Zhong et al. (2024) Y. Zhong, T. Suzuki, H. Liu, K. Liu, Z. Nie, Y. Shi, S. Meng, B. Lv, H. Ding, T. Kanai, J. Itatani, S. Shin, and K. Okazaki, Phys. Rev. Res. 6, 043328 (2024).
- Liu et al. (2025) H. Liu, S. Duan, X. Liu, Z. Liu, S. Wang, L. Gu, J. Huang, W. Yang, J. Liu, D. Qian, Y. Guo, and W. Zhang, Science Bulletin 70, 1211 (2025).
- Fragkos et al. (2025) S. Fragkos, Q. Courtade, O. Tkach, J. Gaudin, D. Descamps, G. Barrette, S. Petit, G. Schönhense, Y. Mairesse, and S. Beaulieu, (2025), 2507.02371 [cond-mat.mtrl-sci] .
- Tkach and Schönhense (2025) O. Tkach and G. Schönhense, Ultramicroscopy 276, 114167 (2025).
- Tkach et al. (2024a) O. Tkach, S. Fragkos, Q. Nguyen, S. Chernov, M. Scholz, N. Wind, S. Babenkov, O. Fedchenko, Y. Lytvynenko, D. Zimmer, A. Hloskovskii, D. Kutnyakhov, F. Pressacco, J. Dilling, L. Bruckmeier, M. Heber, F. Scholz, J. Sobota, J. Koralek, N. Sirica, M. Kallmayer, M. Hoesch, C. Schlueter, L. V. Odnodvorets, Y. Mairesse, K. Rossnagel, H. J. Elmers, S. Beaulieu, and G. Schoenhense, “Multi-mode front lens for momentum microscopy: Part II experiments,” (2024a), arXiv:2401.10084 [cond-mat.mtrl-sci] .
- Moser (2017) S. Moser, Journal of Electron Spectroscopy and Related Phenomena 214, 29 (2017).
- Schusser et al. (2024) J. Schusser, H. Orio, M. Ünzelmann, J. Heßdörfer, M. P. T. Masilamani, F. Diekmann, K. Rossnagel, and F. Reinert, Communications Physics 7, 270 (2024).
- Schusser et al. (2022) J. Schusser, H. Bentmann, M. Ünzelmann, T. Figgemeier, C.-H. Min, S. Moser, J. N. Neu, T. Siegrist, and F. Reinert, Phys. Rev. Lett. 129, 246404 (2022).
- Borisenko et al. (2009) S. V. Borisenko, A. A. Kordyuk, V. B. Zabolotnyy, D. S. Inosov, D. Evtushinsky, B. Büchner, A. N. Yaresko, A. Varykhalov, R. Follath, W. Eberhardt, L. Patthey, and H. Berger, Phys. Rev. Lett. 102, 166402 (2009).
- Huber et al. (2022) M. Huber, Y. Lin, N. Dale, R. Sailus, S. Tongay, R. A. Kaindl, and A. Lanzara, Scientific Reports 12, 15860 (2022).
- Boschini et al. (2024) F. Boschini, M. Zonno, and A. Damascelli, Rev. Mod. Phys. 96, 015003 (2024).
- Buchberger et al. (2025) S. Buchberger, Y. in ’t Veld, A. Rajan, P. A. E. Murgatroyd, B. Edwards, B. K. Saika, N. Kushwaha, M. H. Visscher, J. Berges, D. Carbone, J. Osiecki, C. Polley, T. Wehling, and P. D. C. King, “Persistence of charge ordering instability to coulomb engineering in the excitonic insulator candidate tise2,” (2025), arXiv:2506.01470 [cond-mat.mtrl-sci] .
- Kurtz et al. (2024) F. Kurtz, T. N. Dauwe, S. V. Yalunin, G. Storeck, J. G. Horstmann, H. Böckmann, and C. Ropers, Nature Materials 23, 890 (2024).
- Giustino (2017) F. Giustino, Rev. Mod. Phys. 89, 015003 (2017).
- Lee et al. (2023) H. Lee, S. Poncé, K. Bushick, S. Hajinazar, J. Lafuente-Bartolome, J. Leveillee, C. Lian, J.-M. Lihm, F. Macheda, H. Mori, H. Paudyal, W. H. Sio, S. Tiwari, M. Zacharias, X. Zhang, N. Bonini, E. Kioupakis, E. R. Margine, and F. Giustino, npj Computational Materials 9, 156 (2023).
- Baroni et al. (2001) S. Baroni, S. de Gironcoli, A. Dal Corso, and P. Giannozzi, Rev. Mod. Phys. 73, 515 (2001).
- Giannozzi et al. (2017) P. Giannozzi, O. Andreussi, T. Brumme, O. Bunau, M. Buongiorno Nardelli, M. Calandra, R. Car, C. Cavazzoni, D. Ceresoli, M. Cococcioni, N. Colonna, I. Carnimeo, A. Dal Corso, S. de Gironcoli, P. Delugas, R. A. DiStasio, A. Ferretti, A. Floris, G. Fratesi, G. Fugallo, R. Gebauer, U. Gerstmann, F. Giustino, T. Gorni, J. Jia, M. Kawamura, H.-Y. Ko, A. Kokalj, E. Küçükbenli, M. Lazzeri, M. Marsili, N. Marzari, F. Mauri, N. L. Nguyen, H.-V. Nguyen, A. Otero-de-la Roza, L. Paulatto, S. Poncé, D. Rocca, R. Sabatini, B. Santra, M. Schlipf, A. P. Seitsonen, A. Smogunov, I. Timrov, T. Thonhauser, P. Umari, N. Vast, X. Wu, and S. Baroni, Journal of Physics: Condensed Matter 29, 465901 (2017).
- Sadovskii (2001) M. V. Sadovskii, Physics-Uspekhi 44, 515 (2001).
- Lee et al. (1973) P. A. Lee, T. M. Rice, and P. W. Anderson, Phys. Rev. Lett. 31, 462 (1973).
- McKenzie (1995) R. H. McKenzie, Phys. Rev. B 52, 16428 (1995).
- Pines (1997) D. Pines, Physica C: Superconductivity 282, 273 (1997).
- Kuchinskii et al. (2012) E. Z. Kuchinskii, I. A. Nekrasov, and M. V. Sadovskii, Journal of Experimental and Theoretical Physics 114, 671 (2012).
- Yoshiyama et al. (1986) H. Yoshiyama, Y. Takaoka, N. Suzuki, and K. Motizuki, Journal of Physics C: Solid State Physics 19, 5591 (1986).
- Velebit et al. (2016) K. Velebit, P. Popčević, I. Batistić, M. Eichler, H. Berger, L. Forró, M. Dressel, N. Barišić, and E. Tutiš, Phys. Rev. B 94, 075105 (2016).
- Monney et al. (2015) G. Monney, C. Monney, B. Hildebrand, P. Aebi, and H. Beck, Phys. Rev. Lett. 114, 086402 (2015).
- Lin et al. (2022) Z. Lin, C. Wang, A. Balassis, J. P. Echeverry, A. S. Vasenko, V. M. Silkin, E. V. Chulkov, Y. Shi, J. Zhang, J. Guo, and X. Zhu, Phys. Rev. Lett. 129, 187601 (2022).
- Comby et al. (2022) A. Comby, D. Rajak, D. Descamps, S. Petit, V. Blanchet, Y. Mairesse, J. Gaudin, and S. Beaulieu, Journal of Optics 24, 084003 (2022).
- Medjanik et al. (2017) K. Medjanik, O. Fedchenko, S. Chernov, D. Kutnyakhov, M. Ellguth, A. Oelsner, B. Schönhense, T. R. F. Peixoto, P. Lutz, C.-H. Min, F. Reinert, S. Däster, Y. Acremann, J. Viefhaus, W. Wurth, H. J. Elmers, and G. Schönhense, Nature Materials 16, 615 (2017).
- Tkach et al. (2024b) O. Tkach, S. Fragkos, Q. Nguyen, S. Chernov, M. Scholz, N. Wind, S. Babenkov, O. Fedchenko, Y. Lytvynenko, D. Zimmer, A. Hloskovskii, D. Kutnyakhov, F. Pressacco, J. Dilling, L. Bruckmeier, M. Heber, F. Scholz, J. Sobota, J. Koralek, N. Sirica, M. Kallmayer, M. Hoesch, C. Schlueter, L. V. Odnodvorets, Y. Mairesse, K. Rossnagel, H. J. Elmers, S. Beaulieu, and G. Schoenhense, arXiv , 2408.10104 and 2401.10084 (2024b).
- Xian et al. (2020) R. P. Xian, Y. Acremann, S. Y. Agustsson, M. Dendzik, K. Bühlmann, D. Curcio, D. Kutnyakhov, F. Pressacco, M. Heber, S. Dong, T. Pincelli, J. Demsar, W. Wurth, P. Hofmann, M. Wolf, M. Scheidgen, L. Rettig, and R. Ernstorfer, Scientific Data 7, 442 (2020).
- Xian et al. (2019) R. P. Xian, L. Rettig, and R. Ernstorfer, Ultramicroscopy 202, 133 (2019).
- Hamann (2013) D. R. Hamann, Phys. Rev. B 88, 085117 (2013).
- Perdew et al. (1996) J. P. Perdew, K. Burke, and M. Ernzerhof, Phys. Rev. Lett. 77, 3865 (1996).
- van Setten et al. (2018) M. van Setten, M. Giantomassi, E. Bousquet, M. Verstraete, D. Hamann, X. Gonze, and G.-M. Rignanese, Computer Physics Communications 226, 39 (2018).
- Marzari et al. (2012) N. Marzari, A. A. Mostofi, J. R. Yates, I. Souza, and D. Vanderbilt, Rev. Mod. Phys. 84, 1419 (2012).
- Zhou et al. (2020) J. S. Zhou, L. Monacelli, R. Bianco, I. Errea, F. Mauri, and M. Calandra, Nano Letters 20, 4809 (2020).

### v3

@@ Acknowledgements.
We thank Claude Monney for useful discussions. We thank Nikita Fedorov, Romain Delos, Pierre Hericourt, Rodrigue Bouillaud, Laurent Merzeau, and Frank Blais for technical assistance and Aymen Mahmoudi, Joel Morf, and Frederic Chassot for their assistance during STM measurements. J.S., S.S., and J.M. would like to thank the QM4ST project with Reg. No. CZ.02.01.01/00/22_008/0004572, cofunded by the ERDF as part of the MŠMT. The research leading to these results has received funding from LASERLAB-EUROPE (grant agreement no. 871124, European Union’s Horizon 2020 research and innovation programme). We acknowledge the financial support of the IdEx University of Bordeaux / Grand Research Program "GPR LIGHT". We acknowledge support from ERC Starting Grant ERC-2022-STG No.101076639, Quantum Matter Bordeaux, AAP CNRS Tremplin, and AAP SMR from Université de Bordeaux. This work is part of the ULTRAFAST and TORNADO projects of PEPR LUMA and was supported by the French National Research Agency, as a part of the France 2030 program, under grants ANR-23-EXLU-0002 and ANR-23-EXLU-0004. SF acknowledges funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie 2024 Postdoctoral Fellowship No 101198277 (TopQMat). JS acknowledges funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska‑Curie grant agreement No 101209345—ART.QM funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or European Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work was funded by the Würzburg-Dresden Cluster of Excellence on Complexity, Topology and Dynamics in Quantum Matter – ctd.qmat (EXC 2147, project-id 390858490) and by the DFG through SFB1170 “Tocotronics” and directly via RE 1469/13-2. N.G.E., and D.N. acknowledge financial support from the project "Podizanje znanstvene izvrsnosti Centra za napredne laserske tehnike (CALTboost)" financed by the European Union through the National Recovery and Resilience Plan 2021-2026 (NRPP), European Regional Development Fund for the project ‘Materials for clean energy, advanced sensors and quantum technologies’ (Grant No. PK.1.1.10.0002), and Croatian Science Foundation (Grant no. IP-2025-02-5926). QC also acknowledges support from the TERAQUANTUM project of the Région Nouvelle-Aquitaine. This work was supported by the TWISTnSHINE , funded as project No. LL2314 by Programme ERC CZ.Author contributions
S.F., H.O., S.B., and J.S. performed the experiments, supported by S.S., A.J., Q.C., Y.M., and M.P.T.M.. S.F., H.O., S.B., and J.S. analyzed the experimental data. D.D. and S.P. participated in maintaining the laser system. N.G.E. and D.N. performed the theoretical calculations. F.D. and K.R. synthesized and characterized the samples. B.H. and J.S. performed the STM measurements. S.F., H.O., N.G.E., A.J., S.S., Q.C., M.P.T.M., M.Ü., F.D., B.H., D.D., S.P., F.B., J.M., Y.M., F.R., K.R., D.N., S.B., J.S. contributed to the interpretation and discussion of the results. J.S., S.B., D.N., H.O., and S.F. wrote the manuscript with inputs from coauthors. J.S., S.B., and D.N. conceived and planned the project.
Competing interests
The authors declare no competing interests.